OpenStudy (anonymous):

A hint please? Please help!! (sinx+sin3x)/(2sin2x) = cosx

hartnn (hartnn):

you know the sin C+sin D formula ?

OpenStudy (anonymous):

um i know the formulas sin(2x) = 2 sinxcosx and I know I have to split the Sin(3x) into sin(2x+x) and cos(2x)= 2cos^2x-1 so far I've got sinx + 2sinxcosx(cosx) + 2cos^2x-1(sinx) I don't know if that's right

hartnn (hartnn):

try this : \(\sin C +\sin D = 2 \sin[ (C +D)/2] \cos [(C -D)/2]\)

OpenStudy (anonymous):

lol Yea i saw that on Youtube but i'm just hesitant to use that because my professor didn't teach that formula haha

hartnn (hartnn):

there are multiple ways, this one is short...but if you want to try your method, we can do that.... ?

OpenStudy (anonymous):

yes please! i'm just not sure if what I'm doing is correct. my first step was sinx + sin(2x +x)/ 2sin2x then the top would be sinx + sin(2x)cos(x) + cos(2x)sinx then cos(2x) would change to 2cos^2x-1 sooo it would be sinx + sin(2x)cos(x) + 2cos^2x-1(sinx) am i on the right track?

hartnn (hartnn):

don't miss the brackets! or else next step will become incorrect.. sinx + sin(2x)cos(x) + (2cos^2x-1)(sinx)

hartnn (hartnn):

open the brackets and notice anything getting cancelled ?

OpenStudy (anonymous):

urmm no haha i'm not sure O.o

hartnn (hartnn):

open up the brackets! you can do that, right ?

OpenStudy (anonymous):

do the sin's cancel out?

hartnn (hartnn):

yes!

hartnn (hartnn):

then ?

hartnn (hartnn):

we are very close to the answer :)

OpenStudy (anonymous):

would i be left with sin(2x)cosx + 2cos^2sinx ?

hartnn (hartnn):

ofcourse factor out cos x from that ans see the magic ;)

OpenStudy (anonymous):

cos(sin(2x) + 2cosxsinx) ?

hartnn (hartnn):

and 2 sin x cos x = .. ?

OpenStudy (anonymous):

sin2x!

hartnn (hartnn):

i think you can finish :)

OpenStudy (anonymous):

cos(2sin(2x)/(2sin(2x) = cos!!!!!!! YASSSS OMGGG!!!! :-) XD XD XD :-)